# Positivstellensätze for polynomial matrices with universal quantifiers

### v2

@@ 3. Positivstellensätze for polynomial matrices with UQs under the Archimedean condition
In this section, we shall present a Scherer-Hol-type Positivstellensatz for polynomial matrices with UQs, assuming the Archimedean condition, along with a sparse version in the presence of correlative sparsity.
3.1. Scherer-Hol-type Positivestellensatz
@@
$\mathcal{Q}^{p}(G,\nu)\coloneqq\left\{\Sigma_{0}+\int_{\mathcal{Y}}\left\langle\Sigma,G\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\ \middle|\ \Sigma_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p},\Sigma\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}\,\,\text{are SOS matrices}\right\}.$
Let us first derive the Scherer-Hol-type Positivestellensatz to provide SOS-structured characterizations for polynomial matrices that are PD over $\mathcal{X}$.
@@
Lemma 3.1.
Under Assumption 3.1, the identity matrix $I_{p}$ belongs to the algebraic interior of the cone $\mathcal{Q}^{p}(G,\nu)$ in ${\mathbb{S}}[{\boldsymbol{x}}]^{p}$, i.e., for any $H({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, there exists a real number $t_{0}\geq 0$ such that $H({\boldsymbol{x}})+tI_{p}\in\mathcal{Q}^{p}(G,\nu)$ for all $t\geq t_{0}$.
Proof.
Without loss of generality, we assume that $H({\boldsymbol{x}})=\pm{\boldsymbol{v}}{\boldsymbol{v}}^{\intercal}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}$ for some ${\boldsymbol{v}}\in{\mathbb{R}}^{p}$ and ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}$. We first show by induction that for any ${\boldsymbol{\beta}}\in\mathbb{N}^{n}$, $C^{|{\boldsymbol{\beta}}|}-{\boldsymbol{x}}^{2{\boldsymbol{\beta}}}\in\mathcal{Q}^{1}(G,\nu)$. Note that $C-x_{j}^{2}\in\mathcal{Q}^{1}(G,\nu)$ since
$C-x_{j}^{2}=(C-\|{\boldsymbol{x}}\|^{2})+\sum_{i\neq j}x_{j}^{2}\in\mathcal{Q}^{1}(G,\nu).$
Assume that the claim holds for ${\boldsymbol{\beta}}$ and consider the monomial ${\boldsymbol{x}}^{{\boldsymbol{\beta}}}x_{j_{0}}$ for some $j_{0}\in[n]$. We have
$C^{|{\boldsymbol{\beta}}|+1}-{\boldsymbol{x}}^{2{\boldsymbol{\beta}}}x_{j_{0}}^{2}=C(C^{|{\boldsymbol{\beta}}|}-{\boldsymbol{x}}^{2{\boldsymbol{\beta}}})+{\boldsymbol{x}}^{2{\boldsymbol{\beta}}}(C-x_{j_{0}}^{2})\in\mathcal{Q}^{1}(G,\nu),$
by induction.
Now note that
±$\pm 2{\boldsymbol{v}}{\boldsymbol{v}}^{\intercal}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}={\boldsymbol{v}}{\boldsymbol{v}}^{\intercal}(({\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\pm 1)^{2}-{\boldsymbol{x}}^{2{\boldsymbol{\alpha}}}-1)$
and
${\boldsymbol{v}}{\boldsymbol{v}}^{\intercal}(-{\boldsymbol{x}}^{2{\boldsymbol{\alpha}}}-1)={\boldsymbol{v}}{\boldsymbol{v}}^{\intercal}(C^{|{\boldsymbol{\alpha}}|}-{\boldsymbol{x}}^{2{\boldsymbol{\alpha}}})+(C^{|{\boldsymbol{\alpha}}|}+1)({\boldsymbol{v}}^{\intercal}{\boldsymbol{v}}I_{p}-{\boldsymbol{v}}{\boldsymbol{v}}^{\intercal})-(C^{|{\boldsymbol{\alpha}}|}+1){\boldsymbol{v}}^{\intercal}{\boldsymbol{v}}I_{p}.$
By Cauchy’s inequality, ${\boldsymbol{v}}^{\intercal}{\boldsymbol{v}}I_{p}-{\boldsymbol{v}}{\boldsymbol{v}}^{\intercal}\succeq 0$. Hence, by the above claim,
${\boldsymbol{v}}{\boldsymbol{v}}^{\intercal}(C^{|{\boldsymbol{\alpha}}|}-{\boldsymbol{x}}^{2{\boldsymbol{\alpha}}})+(C^{|{\boldsymbol{\alpha}}|}+1)({\boldsymbol{v}}^{\intercal}{\boldsymbol{v}}I_{p}-{\boldsymbol{v}}{\boldsymbol{v}}^{\intercal})\in\mathcal{Q}^{p}(G,\nu).$
Thus, by letting $t_{0}=\frac{(C^{|{\boldsymbol{\alpha}}|}+1){\boldsymbol{v}}^{\intercal}{\boldsymbol{v}}}{2}$, we have $H({\boldsymbol{x}})+tI_{p}\in\mathcal{Q}^{p}(G,\nu)$ for all $t\geq t_{0}$. ∎
Corollary 3.1.
If $F({\boldsymbol{x}})\not\in\mathcal{Q}^{p}(G,\nu)$, then there exists a linear functional $\mathscr{L}:{\mathbb{S}}[{\boldsymbol{x}}]^{p}\to{\mathbb{R}}$ such that $\mathscr{L}(\mathcal{Q}^{p}(G,\nu))\subseteq{\mathbb{R}}_{+}$, $\mathscr{L}(F)\leq 0$, and $\mathscr{L}(I_{p})=1$.
Proof.
As the cone $\mathcal{Q}^{p}(G,\nu)$ has non-empty algebraic interior by Lemma 3.1, this is due to the Hahn-Banach separation theorem [1, Corollary III.1.7]. ∎
@@
The Borel measure $\nu$ satisfies the multivariable Carleman condition
@@
[18, Proposition 3.2] Suppose that $\nu$ satisfies Assumption 3.2. Then, SOS polynomials are dense in the cone of nonnegative functions in $L^{2}({\mathbb{R}}^{m},\nu)$.
Now, we present our main fundamental result concerning a Scherer-Hol-type Positivstellensatz for polynomial matrices with UQs.
@@
The proof is by contradiction, using the functional $\mathscr{L}$ from Corollary 3.1, and proceeds as follows:
  1. 1.
    Construct a Hilbert space $\mathcal{H}$ and operators $\widehat{\mathbf{A}}\coloneqq(\widehat{A}_{i})_{i\in[n]}$ with $\widehat{A}_{i}\in\mathcal{B}(\mathcal{H})$, where $\mathcal{B}(\mathcal{H})$ denotes the set of bounded linear operators on $\mathcal{H}$. Then construct a state $\Pi$ on the the $p\times p$ matrix algebra $\mathcal{M}_{p}(\mathcal{B}(\mathcal{H}))$ over $\mathcal{B}(\mathcal{H})$ such that $\Pi(F(\widehat{\mathbf{A}}))=\mathscr{L}(F)$;
  2. 2.
    By applying the GNS construction to $(\mathcal{M}_{p}(\mathcal{B}(\mathcal{H})),\Pi)$, construct a Hilbert space $\mathcal{K}_{1}$ with inner product $\langle\cdot,\cdot\rangle_{\mathcal{K}_{1}}$ and operators $\mathbf{A}=(A_{i})_{i\in[n]}$ with $A_{i}\in\mathcal{B}(\mathcal{K}_{1})$ such that $\mathscr{L}(F)=\sum_{i,j=1}^{p}\left\langle F_{ij}(\mathbf{A})\eta_{j},\eta_{i}\right\rangle_{\mathcal{K}_{1}}$ for some $\eta_{i}\in\mathcal{K}_{1}$, $i\in[p]$;
  3. 3.
    By applying the spectral theorem to $\mathbf{A}=(A_{i})_{i\in[n]}$, construct a $p\times p$ PSD matrix-valued measure $\Phi$ on the Borel $\sigma$-algebra $\mathfrak{B}({\mathbb{R}}^{n})$ such that $\mathscr{L}(F)=\int_{{\mathbb{R}}^{n}}F({\boldsymbol{x}})\mathrm{d}\Phi({\boldsymbol{x}})$;
  4. 4.
    Show that the support of $\Phi$ is contained in $\mathcal{X}$ and derive a contradiction by Corollary 3.1.
For the sake of readability, the details of the proof is postponed to Section 6.1. ∎
To conclude this part, let us derive a simple corollary of Theorem 3.1, which will be used in Section 4. We need the following lemma.
@@
Suppose that $\Sigma({\boldsymbol{x}},{\boldsymbol{y}})\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ is an SOS matrix in ${\boldsymbol{x}}$ and ${\boldsymbol{y}}$, then $\int_{\mathcal{Y}}\Sigma({\boldsymbol{x}},{\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})$ is an SOS matrix in ${\boldsymbol{x}}$.
@@
As $\Sigma({\boldsymbol{x}},{\boldsymbol{y}})$ is an SOS matrix, there exists an ×$\ell\times q$ polynomial matrix $T({\boldsymbol{x}},{\boldsymbol{y}})$ for some $\ell\in\mathbb{N}$ such that $\Sigma({\boldsymbol{x}},{\boldsymbol{y}})=T({\boldsymbol{x}},{\boldsymbol{y}})^{\intercal}T({\boldsymbol{x}},{\boldsymbol{y}})$. With $d=\deg_{{\boldsymbol{x}}}(T)$, we could write $T({\boldsymbol{x}},{\boldsymbol{y}})$ as
@@
where $Q({\boldsymbol{y}})$ is the vector of coefficient matrices of $T({\boldsymbol{x}},{\boldsymbol{y}})$ (considered as a polynomial matrix in ${\mathbb{R}}[{\boldsymbol{x}}]^{\ell\times q}$) with respect to $[{\boldsymbol{x}}]_{d}$. Hence,
$\int_{\mathcal{Y}}\Sigma({\boldsymbol{x}},{\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})$ $=\int_{\mathcal{Y}}([{\boldsymbol{x}}]_{d}\otimes I_{q})^{\intercal}(Q({\boldsymbol{y}})^{\intercal}Q({\boldsymbol{y}}))([{\boldsymbol{x}}]_{d}\otimes I_{q})\mathrm{d}\nu({\boldsymbol{y}})$
$=([{\boldsymbol{x}}]_{d}\otimes I_{q})^{\intercal}\int_{\mathcal{Y}}Q({\boldsymbol{y}})^{\intercal}Q({\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})([{\boldsymbol{x}}]_{d}\otimes I_{q})$
As $\int_{\mathcal{Y}}Q({\boldsymbol{y}})^{\intercal}Q({\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})$ is PSD, $\int_{\mathcal{Y}}\Sigma({\boldsymbol{x}},{\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})$ is an SOS matrix in ${\boldsymbol{x}}$. ∎
@@
Let $\mathbf{H}\coloneqq\{H_{1},\ldots,H_{t}\}$ where each $H_{j}\in{\mathbb{S}}[{\boldsymbol{x}}]^{r_{j}}$ for some $r_{j}\in\mathbb{N}$. Suppose that Assumption 3.2 holds and there is $C>0$ such that $C-\|{\boldsymbol{x}}\|^{2}\in\mathcal{Q}^{1}(G,\nu)+\mathcal{Q}^{1}(\mathbf{H})$. If $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is PD on $\mathcal{X}\cap\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid H_{j}({\boldsymbol{x}})\succeq 0,\ j\in[t]\}$, then $F({\boldsymbol{x}})\in\mathcal{Q}^{p}(G,\nu)+\mathcal{Q}^{p}(\mathbf{H})$.
@@
It is clear that $F({\boldsymbol{x}})\succ 0$ on $\mathcal{X}\cap\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid H_{j}({\boldsymbol{x}})\succ 0,\ j\in[t]\}$ if and only if $F({\boldsymbol{x}})\succ 0$ on
@@
By Lemmas 2.2 and 3.2, it is easy to see that $\mathcal{Q}^{p}(\widehat{G},\nu)=\mathcal{Q}^{p}(G,\nu)+\mathcal{Q}^{p}(\mathbf{H})$. So the conclusion follows from Theorem 3.1. ∎
@@ 3.2. A sparse Positivstellensatz for polynomial matrices with UQs
$\widehat{\mathcal{X}}\coloneqq\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\succeq 0,\ \forall j\in[s],\ {\boldsymbol{y}}\in\mathcal{Y}\subset{\mathbb{R}}^{m}\},$
@@
    for =$\text{for }\ell=2,\ldots,t,\ \exists k<\ell\quad\text{s.t.}\quad\mathcal{I}_{\ell}\cap\bigcup_{j<\ell}\mathcal{I}_{j}\subseteq\mathcal{I}_{k};$
@@
$\mathcal{Q}^{p}(\mathbf{G}^{\ell},\nu)\coloneqq\left\{\Sigma_{0}+\sum_{j\in\mathcal{J}_{\ell}}\int_{\mathcal{Y}}\left\langle\Sigma_{j},G_{j}\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\ \middle|\ \Sigma_{0}\in{\mathbb{S}}[{\boldsymbol{x}}(\mathcal{I}_{\ell})]^{p},\Sigma_{j}\in{\mathbb{S}}[{\boldsymbol{x}}(\mathcal{I}_{\ell}),{\boldsymbol{y}}]^{pq_{j}},j\in\mathcal{J}_{\ell},\text{ are SOS matrices}\right\}.$
@@
[12, Lemma 3] Let {}$\{\mathcal{I}_{\ell}\}_{\ell\in[t]}$ satisfy the running intersection property. For any $C>0$, if $f=f_{1}+\cdots+f_{t}$ with $f_{\ell}\in{\mathbb{R}}[{\boldsymbol{x}}(\mathcal{I}_{\ell})]$ satisfies $f>0$ on $[-C,C]^{n}$, then $f=h_{1}+\ldots+h_{t}$ for some $h_{\ell}\in{\mathbb{R}}[{\boldsymbol{x}}(\mathcal{I}_{\ell})]$ with $h_{\ell}>0$ on $[-C,C]^{|\mathcal{I}_{\ell}|}$.
@@
on $[-C,C]^{n}$ for all $k\geq\bar{k}$.
The proof of Proposition 3.3 is postponed to Section 6.2. Note that using Proposition 3.3, we can provide a simpler proof of the Scherer-Hol-type Positivestellensatz (Theorem 3.1) assuming that $\mathcal{Y}$ is compact (see Section 6.2).
@@
on $[-\sqrt{C},\sqrt{C}]^{n}$. Note that for each $j\in[s]$,
@@
and each $h_{\ell}>0$ on $[-\sqrt{C},\sqrt{C}]^{|\mathcal{I}_{\ell}|}$. By Putinar’s Positivstellensatz [41], for each $h_{\ell}$, there exist SOS polynomials $\sigma_{\ell,0},\sigma_{\ell,1}\in{\mathbb{R}}[{\boldsymbol{x}}(\mathcal{I}_{\ell})]$ such that
@@
One might wonder whether the result in Theorem 3.2 holds for a polynomial matrix $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ with $p>1$. Indeed, this is not true even in the absence of UQs; see [21] for a counterexample.

### v4

@@ 3. Positivstellensätze for polynomial matrices with UQs under the Archimedean condition
In this section, assuming the Archimedean condition, we shall present a matrix-valued Positivstellensatz incorporating universal quantifiers, along with a sparse version in the presence of correlative sparsity.
3.1. A Positivstellensatze for polynomial matrices with UQs
@@
$\mathcal{Q}^{p}(G,\nu)\coloneqq\left\{S_{0}+\int_{\mathcal{Y}}\left\langle S,G\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\ \middle|\ S_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p},S\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}\,\,\text{are SOS matrices}\right\}.$
Next, we derive a Positivstellensatze for polynomial matrices with UQs, providing SOS-structured characterizations for polynomial matrices that are PD over $\mathcal{X}$.
@@
The Borel measure $\nu$ satisfies the multivariate Carleman condition
@@
[16, Proposition 3.2] Suppose that $\nu$ satisfies Assumption 3.2. Then, SOS polynomials are dense in the cone of nonnegative functions in $L^{2}({\mathbb{R}}^{m},\nu)$.
Proposition 3.2.
Suppose that $\nu$ satisfies Assumption 3.2. Then
$\mathcal{X}=\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid H({\boldsymbol{x}})\succeq 0,\ \forall H\in\mathcal{Q}^{p}(G,\nu)\}.$ (3)
Proof.
By the definition of $\mathcal{Q}^{p}(G,\nu)$, we only need to prove that $\mathcal{X}$ contains the set on the right-hand side of the equation in (3). Fix ${\boldsymbol{u}}\in{\mathbb{R}}^{n}$ with $H({\boldsymbol{u}})\succeq 0$ for all $H\in\mathcal{Q}^{p}(G,\nu)$. Suppose on the contrary that ${\boldsymbol{u}}\not\in\mathcal{X}$, i.e., there exists ${\boldsymbol{w}}\in\mathcal{Y}$ such that $G({\boldsymbol{u}},{\boldsymbol{w}})\not\succeq 0$. Then, there is a ball $\mathcal{O}\subset{\mathbb{R}}^{m}$ with a radius $\rho>0$ around $\boldsymbol{w}$} such that $G({\boldsymbol{u}},{\boldsymbol{y}})\not\succeq 0$ on $2\mathcal{O}$. We may assume that there exists ${\boldsymbol{v}}\in{\mathbb{R}}^{q}$ and $\delta>0$ such that ${\boldsymbol{v}}^{\intercal}G({\boldsymbol{u}},{\boldsymbol{y}}){\boldsymbol{v}}\leq-\delta$ on $2\mathcal{O}$. Define a continuous function $h({\boldsymbol{y}})$ on ${\mathbb{R}}^{m}$ by $h({\boldsymbol{y}})=2\rho-\|{\boldsymbol{y}}-{\boldsymbol{w}}\|$ for ${\boldsymbol{y}}\in 2\mathcal{O}$ and $h({\boldsymbol{y}})=0$ otherwise. By Proposition 3.1, there exists a sequence of SOS polynomials $\{\sigma_{k}\}_{k}$ in ${\mathbb{R}}[{\boldsymbol{y}}]$ that converges to $h$ in the $L^{2}$-norm. Hence,
$\lim_{k\to\infty}\int_{\mathcal{Y}}{\boldsymbol{v}}^{\intercal}G({\boldsymbol{u}},{\boldsymbol{y}}){\boldsymbol{v}}\sigma_{k}({\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})=\int_{\mathcal{Y}}{\boldsymbol{v}}^{\intercal}G({\boldsymbol{u}},{\boldsymbol{y}}){\boldsymbol{v}}h({\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})$
$=$ $\int_{\mathcal{Y}\cap 2\mathcal{O}}{\boldsymbol{v}}^{\intercal}G({\boldsymbol{u}},{\boldsymbol{y}}){\boldsymbol{v}}(2\rho-\|{\boldsymbol{y}}-{\boldsymbol{w}}\|)\mathrm{d}\nu({\boldsymbol{y}})\leq\int_{\mathcal{Y}\cap\mathcal{O}}-\delta\rho\mathrm{d}\nu({\boldsymbol{y}})=-\delta\rho\nu(\mathcal{Y}\cap\mathcal{O})<0,$
where the last inequality is due to the fact that $\mathrm{supp}(\nu)=\mathcal{Y}$ and thus $\nu(\mathcal{Y}\cap\mathcal{O})>0$. Note that for all $k\in\mathbb{N}$,
$\int_{\mathcal{Y}}{\boldsymbol{v}}^{\intercal}G({\boldsymbol{u}},{\boldsymbol{y}}){\boldsymbol{v}}\sigma_{k}({\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})=\frac{1}{p}\mbox{tr}\left(\int_{\mathcal{Y}}\left\langle\sigma_{k}({\boldsymbol{y}})I_{p}\otimes{\boldsymbol{v}}{\boldsymbol{v}}^{\intercal},G({\boldsymbol{u}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\right)\geq 0,$
since $\sigma_{k}({\boldsymbol{y}})I_{p}\otimes{\boldsymbol{v}}{\boldsymbol{v}}^{\intercal}$ is an SOS and $H({\boldsymbol{u}})\succeq 0$ for all $H\in\mathcal{Q}^{p}(G,\nu)$. A contradiction follows. ∎
Now, we present our main result concerning a Positivstellensatz for polynomial matrices with UQs.
@@
Since the quadratic module $\mathcal{Q}^{p}(G,\nu)$ is Archimedean and the equality in (3) holds, the conclusion follows from the fundamental Positivstellensatz for matrix algebras of polynomials [44, Theorem 10.25]. ∎
Next, we derive a corollary of Theorem 3.1, which will be used in Section 4.
@@
Suppose that $S({\boldsymbol{x}},{\boldsymbol{y}})\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ is an SOS matrix in $\boldsymbol{x}$} and $\boldsymbol{y}$}. Then $\int_{\mathcal{Y}}S({\boldsymbol{x}},{\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})$ is an SOS matrix in $\boldsymbol{x}$}.
@@
As $S({\boldsymbol{x}},{\boldsymbol{y}})$ is an SOS matrix, there exists an $\ell\times q$ polynomial matrix $T({\boldsymbol{x}},{\boldsymbol{y}})$ for some $\ell\in\mathbb{N}$ such that $S({\boldsymbol{x}},{\boldsymbol{y}})=T({\boldsymbol{x}},{\boldsymbol{y}})^{\intercal}T({\boldsymbol{x}},{\boldsymbol{y}})$. With $d=\deg_{{\boldsymbol{x}}}T$, we could write $T({\boldsymbol{x}},{\boldsymbol{y}})$ as
@@
where $Q({\boldsymbol{y}})$ is the vector of coefficient matrices of $T({\boldsymbol{x}},{\boldsymbol{y}})$ (considered as a polynomial matrix in ${\mathbb{R}}[{\boldsymbol{x}}]^{\ell\times q}$) with respect to ${\boldsymbol{x}}]_{d}$. Hence,
$\int_{\mathcal{Y}}S({\boldsymbol{x}},{\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})=\int_{\mathcal{Y}}([{\boldsymbol{x}}]_{d}\otimes I_{q})^{\intercal}(Q({\boldsymbol{y}})^{\intercal}Q({\boldsymbol{y}}))([{\boldsymbol{x}}]_{d}\otimes I_{q})\mathrm{d}\nu({\boldsymbol{y}})=([{\boldsymbol{x}}]_{d}\otimes I_{q})^{\intercal}\int_{\mathcal{Y}}Q({\boldsymbol{y}})^{\intercal}Q({\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})([{\boldsymbol{x}}]_{d}\otimes I_{q})$
As $\int_{\mathcal{Y}}Q({\boldsymbol{y}})^{\intercal}Q({\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})$ is PSD, $\int_{\mathcal{Y}}S({\boldsymbol{x}},{\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})$ is an SOS matrix in $\boldsymbol{x}$}. ∎
@@
Let $\mathbf{H}\coloneqq\{H_{1},\ldots,H_{t}\}$ where each $H_{j}\in{\mathbb{S}}[{\boldsymbol{x}}]^{r_{j}}$ for some $r_{j}\in\mathbb{N}$. Suppose that Assumption 3.2 holds and there exists $C>0$ such that $C-\|{\boldsymbol{x}}\|^{2}\in\mathcal{Q}^{1}(G,\nu)+\mathcal{Q}^{1}(\mathbf{H})$. If $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is PD on $\mathcal{X}\cap\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid H_{j}({\boldsymbol{x}})\succeq 0,\ j\in[t]\}$, then $F({\boldsymbol{x}})\in\mathcal{Q}^{p}(G,\nu)+\mathcal{Q}^{p}(\mathbf{H})$.
@@
It is clear that $F({\boldsymbol{x}})\succ 0$ on $\mathcal{X}\cap\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid H_{j}({\boldsymbol{x}})\succeq 0,\ j\in[t]\}$ if and only if $F({\boldsymbol{x}})\succ 0$ on
@@
By Lemmas 2.2 and 3.1, it is easy to see that $\mathcal{Q}^{p}(\widehat{G},\nu)=\mathcal{Q}^{p}(G,\nu)+\mathcal{Q}^{p}(\mathbf{H})$. So, the conclusion follows from Theorem 3.1. ∎
@@ 3.2. A sparse Positivstellensatz for polynomial matrices with UQs
$\widehat{\mathcal{X}}\coloneqq\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\succeq 0,\ \forall j\in[s],\ {\boldsymbol{y}}\in\mathcal{Y}\subset{\mathbb{R}}^{m}\},$ (4)
@@
    $\text{for }\ell=2,\ldots,t,\ \exists k<\ell\quad\text{s.t.}\quad\mathcal{I}_{\ell}\cap\bigcup_{j<\ell}\mathcal{I}_{j}\subseteq\mathcal{I}_{k};$
@@
$\mathcal{Q}^{p}(\mathbf{G}^{\ell},\nu)\coloneqq\left\{S_{0}+\sum_{j\in\mathcal{J}_{\ell}}\int_{\mathcal{Y}}\left\langle S_{j},G_{j}\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\ \middle|\ S_{0}\in{\mathbb{S}}[{\boldsymbol{x}}(\mathcal{I}_{\ell})]^{p},S_{j}\in{\mathbb{S}}[{\boldsymbol{x}}(\mathcal{I}_{\ell}),{\boldsymbol{y}}]^{pq_{j}},j\in\mathcal{J}_{\ell},\text{ are SOS matrices}\right\}.$
@@
[11, Lemma 3] Let $\{\mathcal{I}_{\ell}\}_{\ell\in[t]}$ satisfy the running intersection property. For any $C>0$, if $f=f_{1}+\cdots+f_{t}$ with $f_{\ell}\in{\mathbb{R}}[{\boldsymbol{x}}(\mathcal{I}_{\ell})]$ satisfies $f>0$ on $C,C]^{n}$, then $f=h_{1}+\ldots+h_{t}$ for some $h_{\ell}\in{\mathbb{R}}[{\boldsymbol{x}}(\mathcal{I}_{\ell})]$ with $h_{\ell}>0$ on $C,C]^{|\mathcal{I}_{\ell}|}$.
@@
on $C,C]^{n}$ for all $k\geq\bar{k}$.
The proof of Proposition 3.4 is postponed to Section 6.1. Note that using Proposition 3.4, we can provide a constructive proof of Theorem 3.1 assuming that $\mathcal{Y}$ is compact (see Section 6.1).
@@
on $\sqrt{C},\sqrt{C}]^{n}$. Note that for each $j\in[s]$,
@@
and each $h_{\ell}>0$ on $\sqrt{C},\sqrt{C}]^{|\mathcal{I}_{\ell}|}$. By Putinar’s Positivstellensatz [39], for each $h_{\ell}$, there exist SOS polynomials $\sigma_{\ell,0},\sigma_{\ell,1}\in{\mathbb{R}}[{\boldsymbol{x}}(\mathcal{I}_{\ell})]$ such that
@@
One might wonder whether the result of Theorem 3.2 holds for a polynomial matrix $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ with $p>1$. Indeed, this is not true even in the absence of UQs; see [32] for a counterexample.